# Speech Language Models for Under-Represented Languages: Insights from Wolof

### v1

@@ Abstract
We present our journey in training a speech language model for Wolof, an underrepresented language spoken in West Africa, and share key insights. We first emphasize the importance of collecting large-scale, spontaneous, high-quality speech data, and show that continued pretraining HuBERT on this dataset outperforms both the base model and African-centric models on ASR. We then integrate this speech encoder into a Wolof LLM to train the first Speech LLM for this language, extending its capabilities to tasks such as speech translation. Furthermore, we explore training the Speech LLM to perform multi-step Chain-of-Thought before transcribing or translating. Our results show that the Speech LLM not only improves speech recognition but also performs well in speech translation. The models and the code will be openly shared.
@@ 3.3 Evaluation on an ASR task
Results. Table 2 shows that our model improves over Meta/HuBERT-Base and Orange/HuBERT-Base, despite being trained on far less speech data than the latter. We attribute this superior performance to continued pretraining (Insight 1) and to the Wolof-centric, high-quality data collection (Insight 2). Orange/HuBERT-Base underperforms compared to Meta/HuBERT-Base because the former is trained from scratch and probably with fewer epochs. This is also a reason to do continued pretraining: it requires fewer training steps to perform well.
@@
Table 2: WER results of the continued pretrained HuBERT compared to the initial model (Meta/HuBERT-Base) and an African-centric model Orange/HuBERT-Base. 960 $\rightarrow$ 860 means the initial model is trained on 960 hours and continued pretrained on 860 hours of Wolof speech.
@@ 3.4 Discussion
3.5 Training Wolof LLM
Model. We finetune the Wolof LLM from Qwen2.5 3B [cite]. While other LLMs can also be used, we choose this LLM not only for its good general performance but also for its open license, enabling us to redistribute the fine-tuned model freely.

### v2

@@ Abstract
We present our journey in training a speech language model for Wolof, an underrepresented language spoken in West Africa, and share key insights. We first emphasize the importance of collecting large-scale, spontaneous, high-quality unsupervised speech data, and show that continued pretraining HuBERT on this dataset outperforms both the base model and African-centric models on ASR. We then integrate this speech encoder into a Wolof LLM to train the first Speech LLM for this language, extending its capabilities to tasks such as speech translation. Furthermore, we explore training the Speech LLM to perform multi-step Chain-of-Thought before transcribing or translating. Our results show that the Speech LLM not only improves speech recognition but also performs well in speech translation. The models and the code will be openly shared.
@@ 3.3 Evaluation on an ASR task
Results. Table 2 shows that our model improves over Meta/HuBERT-Base and Orange/HuBERT-Base, despite being trained on far less speech data than the latter. We attribute this superior performance to continued pretraining (Insight 1) and to the Wolof-centric, high-quality data collection (Insight 2). Orange/HuBERT-Base underperforms compared to Meta/HuBERT-Base because the former is trained from scratch and probably with fewer epochs. This is also a reason to conduct continued pretraining: it requires fewer training steps to achieve good performance.
@@
Table 2: WER results of the continued pretrained HuBERT compared to the initial model (Meta/HuBERT-Base) and an African-centric model Orange/HuBERT-Base. 960 $\rightarrow$ 860 means the initial model is trained on 960 hours and continued pretrained on 860 hours of Wolof speech. In any case, the WER remains high because, contrary to English, the orthography of Wolof is not standardized since it’s primarily an oral language.
@@ 4 Speech Language Model
4.1 Training Wolof LLM
Model. We finetune the Wolof LLM from Qwen2.5 3B [cite]. While other LLMs can also be used, we chose this LLM not only for its good general performance but also for its open license, enabling us to redistribute the fine-tuned model freely.